OpenStudy (anonymous):

The curve ax^2+bx+c passes through the point (1,2) and is tangent to the line y=x at the origin. Find a, b, and c.

OpenStudy (ujjwal):

Is it y=ax^2+bx+c ?

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

Thoughts?

OpenStudy (ujjwal):

Substitute the point (1,2) in y=ax^2+bx+c you will get your 1st equation Now the curve is tangent to y=x at (0,0) So, substitute (0,0) in equation y=ax^2+bx+c you will get a second equation the 1st derivative of y=ax^2+bx+c i.e. y'=2ax+b at (0,0) will give you slope of tangent of the curve at (0,0).. Substituting the point (0,0) in y'=2ax+b you get slope(m)=b but the slope of y=x is 1 so, you have b=1 So, you have b=1 and two other equations.. Solve them to get a and c..

OpenStudy (anonymous):

Okay, between this and my friend, I've figured it out! Thank you, sir!